A projectile is fired at an angle 15 Β° from a gun that is 45 m above the flat ground, emerging


from the gun with a speed of 300 m/s.


(a) How long does the projectile remain in air?


(b) At what horizontal distance from the firing ground does it strike the ground.


(c) What is the velocity of the projectile when it hits the ground?

1
Expert's answer
2021-11-16T09:59:17-0500

(a) The time:


"T=\frac{v\sin\theta}{g}+\sqrt{\frac{2[h_0+(v\sin\theta)^2/(2g)]}{g}}=16.4\text{ s}."

(b) The range:


"R=Tv\cos\theta=4752\text{ m}."

(c) The velocity:


"u=\sqrt{[v\cos\theta]^2+\Big[\sqrt{2g(h_0+[(v\sin\theta)^2/(2g)]}\Big]^2}=\\=301\text{ m/s}."